# Improving Variational Autoencoder using Random Fourier Transformation: An Aviation Safety Anomaly Detection Case-Study

### v1

@@ 1 Introduction
Neural networks have emerged as indispensable tools in the realm of artificial intelligence, fueling advancements in image recognition, anomaly detection, and generative modeling, among other domains [de2022review]. Central to their effectiveness is their ability to automatically learn and extract meaningful features from raw data, enabling them to discover complex patterns and relationships. Among the neural architectures, autoencoders stand out for their capacity to diminish dimensionality and capture essential data representations. Autoencoders, first introduced by [rumelhart1985learning], represent a class of neural network architectures designed for unsupervised learning, dimensionality reduction, and feature learning [baldi2012autoencoders]. Comprising both an encoder and a decoder, autoencoders, specifically the compressive variant commonly referred to as autoencoders, seek to map input data to a lower-dimensional representation and then reconstruct the original data from this compressed form. Autoencoders have found applications in various fields, including anomaly detection, denoising, and data compression, through their ability to learn meaningful feature representations.
@@
Variational Autoencoders (VAEs) have emerged as a significant advancement in generative modeling, extending the capabilities of traditional autoencoders. They effectively model complex data distributions and generate novel, high-quality data samples by learning a probabilistic distribution over the latent space [kingma2019introduction, memarzadeh2020unsupervised, akbari2023probabilistic]. Unlike traditional autoencoders, VAEs incorporate a probabilistic latent space, enabling them to capture uncertainty and generate diverse samples. The reparameterization trick facilitates efficient training by sampling from a simpler distribution and transforming the samples using mean and variance parameters output by the encoder. VAEs typically learn continuous latent representations, enabling smooth interpolation between data points in the latent space. During training, VAEs optimize an objective function that balances a reconstruction loss, measuring fidelity to the input, and a regularization term, typically the Kullback Leibler (KL) divergence between the learned latent distribution and a prior distribution. This encourages the latent space to follow a predefined structure, promoting smoothness and continuity. The probabilistic nature, continuous latent space, and generative modeling capabilities position VAEs as an improvement to autoencoders and make them versatile tools for various data-driven applications, including image synthesis, data augmentation, and robust anomaly detection.
@@
Despite the promising capabilities of neural networks, including autoencoders and VAEs, in learning data with complex patterns, they still struggle with capturing the full complexity and subtlety of the underlying data distributions. Specifically, neural networks struggle with learning high-frequency patterns that often correspond to fine-grained details, subtle variations, and localized features within the data [tancik2020fourier, xu2019frequency, luo2019theory, xu2022overview]. Neural networks, including autoencoders and VAEs, often struggle to represent and model these high-frequency patterns effectively, primarily focusing on the more dominant low-frequency components [xu2019frequency, xu2022overview, luo2019theory]. This behavior of neural networks referred to as ”spectral bias”, calls for a better understanding of neural network behavior and how it can be mitigated to increase the effectiveness and efficiency of these models.
@@
Numerous research efforts have been devoted to addressing the high-frequency learning difficulty encountered by neural networks [he2016deep, xu2019frequency, tancik2020fourier]. he2016deep proposed Residual Networks (ResNets) which mitigate the spectral bias by learning the residuals (i.e. high-pass frequencies) instead of the signal in each block. Multiple strategies have been proposed and successfully applied to implicit models, offering a more robust positional encoding technique [zheng2021rethinking, damodaran2023improved] that enhances the model performance. tancik2020fourier proposed an integration of sinusoidal mapping techniques applied to the input data. Using the neural tangent kernel (NTK) [jacot2018neural], the authors illustrated that simple neural networks are impractically slow in learning high-frequency signals in data, and demonstrated significant promise of Random Fourier Transformation in the low-dimensional domain.
@@
Random Fourier Transformation (RFT) is a versatile technique that leverages principles from Fourier analysis to transform input data into a low-dimensional Euclidean inner product feature space [rahimi2007random]. By introducing the concept of randomness into this transformation, RFT offers a computationally efficient and scalable means of approximating kernel functions. It does so by using random feature maps, thereby enabling the efficient application of kernel methods without explicitly computing kernel matrices. RFT projects the data into a unit circle space where the data’s inner product is an unbiased estimator of the kernel. RFT has gained prominence in machine learning, particularly in scenarios where traditional kernel methods, such as the kernel trick in Support Vector Machines, are computationally infeasible due to their high time and memory complexity [rahimi2007random]. RFTs approximate shift-invariant kernels $k(x,x^{\prime})=k(x-x^{\prime})$ which can be viewed as an augmentation technique for shift-prune neural networks architecture (i.e. Dense layers, recurrent layers) to improve their robustness and generalization [vaish2024fourier, xu2023fourier, rahimi2007random]. The impact of RFTs in convolutional blocks is diminished due to convolutional and pooling layers’ inherently shift-invariant characteristics. Furthermore, Random Fourier Transformations closely follow the harmonic function concept if all the requirements of orthogonality are respected by the initial randomization [harris1978use, stein1971introduction] which can be regarded as regularization term for neural networks that helps them in the training process.
@@
By further investigating the architecture of RFT, as a precursor layer in neural networks, it becomes obvious that the layer can be thought of as two shared weight random and fixed convolutional layers with a kernel size of 1 that expands the channels of the input data into the pre-defined number of random Fourier channels in parallel with sine and cosine activation functions followed by a concatenation operator. The harmonic activation functions used in this form are also reported to improve the performance of implicit neural networks [sitzmann2020implicit] and neural networks, in general, [liang2021reproducing]. Given the new form of RFT, we described to generate the unbiased kernel, we further explored the possibility of training the kernel for maximum optimization.
@@ 2.1 Autoencoder
Autoencoders, introduced by [rumelhart1985learning], consists of two main components: (1) encoder $(g_{\phi}(.))$: which encodes the data ($x$) into latent representation ($z$), and (2) decoder $(f_{\theta}(.))$: which transforms latent space back to the input data space 1. Autoencoders are mainly categorized into expansive and compressive types, based on larger or smaller latent representations compared to the original data size. Compressive autoencoders are the dominant type due to the constant need for dimensionality reduction, especially while dealing with high-dimensional data. Compressive autoencoders are also considered a form of regularization to prevent ”copy-the-input” artifacts when the network learns to copy the input instead of learning its underlying features.
@@ 2.2 Variational Autoencoder
The latent space in autoencoders is deterministic, representing a fixed transformation of the input data. However, as the need to capture the inherent variability within data became evident, the transition to VAEs emerged. This transition introduces the concept of probabilistic modeling and variational inference into the autoencoder framework [kingma2013auto]. VAEs redefine the latent space as a probability distribution, enabling the encoding process to capture the stochastic nature of data. By incorporating the KL divergence term in their loss function, VAEs strike a balance between data reconstruction and regularization, thereby accommodating encodings with probabilistic variability.
Variational Inference is the cornerstone for VAEs, allowing them to approximate complex probability distributions in a computationally efficient manner. Variational Inference operates by framing the estimation of a target distribution as an optimization problem, seeking an approximate distribution that best aligns with the true posterior [kingma2013auto, kingma2019introduction]. The proposed training process sidesteps the often intractable direct computation of posterior distributions, a formidable challenge in probabilistic modeling. Consequently, Variational Inference plays a pivotal role in probabilistic graphical models (e.g. VAEs) and enables probabilistic learning of complex data in expressive latent configurations. Mathematically, the goal of Variational Inference is to find the best approximation, often denoted as $q(\mathbf{z})$, to the true posterior distribution $p(\mathbf{z}|\mathbf{x})$, where $\mathbf{z}$ represents latent variables and $\mathbf{x}$ are the observed data. This approximation is achieved by minimizing the KL divergence between the true posterior and the approximate distribution:
@@ 3.1 Low Dimensional Datasets
The benchmarks used in this study are adapted from [xu2019frequency] to illustrate how neural networks perform when trained on straightforward datasets. These benchmarks serve as a means to evaluate the behavior and capabilities of neural networks when faced with simplified data structures. By leveraging these benchmarks, we aim to gain insights into how neural networks handle and learn from basic datasets, providing a foundational understanding of their performance in more complex scenarios. The utilization of established benchmarks also allows for comparisons with existing research, facilitating a deeper analysis of neural network behavior across different experimental setups and dataset complexities.
@@ 3.2 High Dimensional Datasets
The dataset used for this study focuses on detecting anomalies related to the deployment of flaps and path deviations during the final approach to landing in commercial aviation [memarzadeh2022multiclass]. The anomalies in the dataset are characterized by a delay in flap deployment, lasting approximately 160 seconds, as identified by subject matter experts. It’s important to note that while this labeled anomaly is included in the dataset, it is not utilized as an input during unsupervised training. The dataset consists of ten time-series representing aeronautical sensor outputs, encompassing parameters such as aircraft position, orientation, speed, and the positions of control surfaces.
@@ 4.1 Benchmark Experiments
In our exploration of model training dynamics, we investigate the Frequency-Principle analysis as proposed by xu2019frequency [xu2019frequency]. This analysis method, detailed in their work, involves decomposing signals into Fourier frequencies. By employing this approach, we aim to gain deeper insights into the spectral bias phenomenon prevalent in neural network learning processes. Spectral bias refers to the tendency of neural networks to favor learning certain frequency components of input data over others, which can significantly influence model performance and generalization.
xu2019frequency’s [xu2019frequency] Frequency-Principle analysis serves as a valuable framework for dissecting how neural networks interact with and prioritize different frequency components during training. Through this analysis, we seek to uncover patterns and tendencies in model behavior related to frequency-specific learning preferences. By understanding these dynamics, we can make informed adjustments to training strategies, network architectures, or preprocessing techniques to mitigate spectral bias effects and enhance overall model robustness and performance across various datasets and tasks.
@@ 5 Conclusion
6 Appendix
Figure 11: Detailed model architecture for CVAE and CAE models. The primary distinction between CVAE and CAE lies in their respective methods of calculating the ”Latent Space.” CVAE achieves this by defining multivariate Gaussian distribution parameters ($\mu$ and $\sigma$), followed by sampling from the distribution. Conversely, CAE directly computes the latent representation.

### v3

@@ 1 Introduction
Neural networks have emerged as indispensable tools in the realm of artificial intelligence, fueling advancements in image recognition, anomaly detection, and generative modeling, among other domains [4]. Central to their effectiveness is their ability to automatically learn and extract meaningful features from raw data, enabling them to discover complex patterns and relationships. Among the neural architectures, autoencoders stand out for their capacity to diminish dimensionality and capture essential data representations. Autoencoders, first introduced by [15], represent a class of neural network architectures designed for unsupervised learning, dimensionality reduction, and feature learning [2]. Comprising both an encoder and a decoder, autoencoders, specifically the compressive variant commonly referred to as autoencoders, seek to map input data to a lower-dimensional representation and then reconstruct the original data from this compressed form. Autoencoders have found applications in various fields, including anomaly detection, denoising, and data compression, through their ability to learn meaningful feature representations.
@@
Variational Autoencoders (VAEs) have emerged as a significant advancement in generative modeling, extending the capabilities of traditional autoencoders. They effectively model complex data distributions and generate novel, high-quality data samples by learning a probabilistic distribution over the latent space [8, 12, 1]. Unlike traditional autoencoders, VAEs incorporate a probabilistic latent space, enabling them to capture uncertainty and generate diverse samples. The reparameterization trick facilitates efficient training by sampling from a simpler distribution and transforming the samples using mean and variance parameters output by the encoder. VAEs typically learn continuous latent representations, enabling smooth interpolation between data points in the latent space. During training, VAEs optimize an objective function that balances a reconstruction loss, measuring fidelity to the input, and a regularization term, typically the Kullback Leibler (KL) divergence between the learned latent distribution and a prior distribution. This encourages the latent space to follow a predefined structure, promoting smoothness and continuity. The probabilistic nature, continuous latent space, and generative modeling capabilities position VAEs as an improvement to autoencoders and make them versatile tools for various data-driven applications, including image synthesis, data augmentation, and robust anomaly detection.
@@
Despite the promising capabilities of neural networks, including autoencoders and VAEs, in learning data with complex patterns, they still struggle with capturing the full complexity and subtlety of the underlying data distributions. Specifically, neural networks struggle with learning high-frequency patterns that often correspond to fine-grained details, subtle variations, and localized features within the data [18, 21, 11, 22]. Neural networks, including autoencoders and VAEs, often struggle to represent and model these high-frequency patterns effectively, primarily focusing on the more dominant low-frequency components [21, 22, 11]. This behavior of neural networks referred to as ”spectral bias”, calls for a better understanding of neural network behavior and how it can be mitigated to increase the effectiveness and efficiency of these models.
@@
Numerous research efforts have been devoted to addressing the high-frequency learning difficulty encountered by neural networks [6, 21, 18]. He et al. [6] proposed Residual Networks (ResNets) which mitigate the spectral bias by learning the residuals (i.e. high-pass frequencies) instead of the signal in each block. Multiple strategies have been proposed and successfully applied to implicit models, offering a more robust positional encoding technique [23, 3] that enhances the model performance. Tancik et al. [18] proposed an integration of sinusoidal mapping techniques applied to the input data. Using the neural tangent kernel (NTK) [7], the authors illustrated that simple neural networks are impractically slow in learning high-frequency signals in data, and demonstrated significant promise of Random Fourier Transformation in the low-dimensional domain.
@@
Random Fourier Transformation (RFT) is a versatile technique that leverages principles from Fourier analysis to transform input data into a low-dimensional Euclidean inner product feature space [14]. By introducing the concept of randomness into this transformation, RFT offers a computationally efficient and scalable means of approximating kernel functions. It does so by using random feature maps, thereby enabling the efficient application of kernel methods without explicitly computing kernel matrices. RFT projects the data into a unit circle space where the data’s inner product is an unbiased estimator of the kernel. RFT has gained prominence in machine learning, particularly in scenarios where traditional kernel methods, such as the kernel trick in Support Vector Machines, are computationally infeasible due to their high time and memory complexity [14]. RFTs approximate shift-invariant kernels $k(x,x^{\prime})=k(x-x^{\prime})$ which can be viewed as an augmentation technique for shift-prune neural networks architecture (i.e. Dense layers, recurrent layers) to improve their robustness and generalization [19, 20, 14]. The impact of RFTs in convolutional blocks is diminished due to convolutional and pooling layers’ inherently shift-invariant characteristics. Furthermore, Random Fourier Transformations closely follow the harmonic function concept if all the requirements of orthogonality are respected by the initial randomization [5, 17] which can be regarded as regularization term for neural networks that helps them in the training process.
@@
By further investigating the architecture of RFT, as a precursor layer in neural networks, it becomes obvious that the layer can be thought of as two shared weight random and fixed convolutional layers with a kernel size of 1 that expands the channels of the input data into the pre-defined number of random Fourier channels in parallel with sine and cosine activation functions followed by a concatenation operator. The harmonic activation functions used in this form are also reported to improve the performance of implicit neural networks [16] and neural networks, in general, [10]. Given the new form of RFT, we described to generate the unbiased kernel, we further explored the possibility of training the kernel for maximum optimization.
@@ 2.1 Autoencoder
Autoencoders, introduced by [15], consists of two main components: (1) encoder $(g_{\phi}(.))$: which encodes the data ($x$) into latent representation ($z$), and (2) decoder $(f_{\theta}(.))$: which transforms latent space back to the input data space 1. Autoencoders are mainly categorized into expansive and compressive types, based on larger or smaller latent representations compared to the original data size. Compressive autoencoders are the dominant type due to the constant need for dimensionality reduction, especially while dealing with high-dimensional data. Compressive autoencoders are also considered a form of regularization to prevent ”copy-the-input” artifacts when the network learns to copy the input instead of learning its underlying features.
@@ 2.2 Variational Autoencoder
The latent space in autoencoders is deterministic, representing a fixed transformation of the input data. However, as the need to capture the inherent variability within data became evident, the transition to VAEs emerged. This transition introduces the concept of probabilistic modeling and variational inference into the autoencoder framework [9]. VAEs redefine the latent space as a probability distribution, enabling the encoding process to capture the stochastic nature of data. By incorporating the KL divergence term in their loss function, VAEs strike a balance between data reconstruction and regularization, thereby accommodating encodings with probabilistic variability.
Variational Inference is the cornerstone for VAEs, allowing them to approximate complex probability distributions in a computationally efficient manner. Variational Inference operates by framing the estimation of a target distribution as an optimization problem, seeking an approximate distribution that best aligns with the true posterior [9, 8]. The proposed training process sidesteps the often intractable direct computation of posterior distributions, a formidable challenge in probabilistic modeling. Consequently, Variational Inference plays a pivotal role in probabilistic graphical models (e.g. VAEs) and enables probabilistic learning of complex data in expressive latent configurations. Mathematically, the goal of Variational Inference is to find the best approximation, often denoted as $q(\mathbf{z})$, to the true posterior distribution $p(\mathbf{z}|\mathbf{x})$, where $\mathbf{z}$ represents latent variables and $\mathbf{x}$ are the observed data. This approximation is achieved by minimizing the KL divergence between the true posterior and the approximate distribution:
@@ 3.1 Low Dimensional Datasets
The benchmarks used in this study are adapted from [21] to illustrate how neural networks perform when trained on straightforward datasets. These benchmarks serve as a means to evaluate the behavior and capabilities of neural networks when faced with simplified data structures. By leveraging these benchmarks, we aim to gain insights into how neural networks handle and learn from basic datasets, providing a foundational understanding of their performance in more complex scenarios. The utilization of established benchmarks also allows for comparisons with existing research, facilitating a deeper analysis of neural network behavior across different experimental setups and dataset complexities.
@@ 3.2 High Dimensional Datasets
The dataset used for this study focuses on detecting anomalies related to the deployment of flaps and path deviations during the final approach to landing in commercial aviation [13]. The anomalies in the dataset are characterized by a delay in flap deployment, lasting approximately 160 seconds, as identified by subject matter experts. It’s important to note that while this labeled anomaly is included in the dataset, it is not utilized as an input during unsupervised training. The dataset consists of ten time-series representing aeronautical sensor outputs, encompassing parameters such as aircraft position, orientation, speed, and the positions of control surfaces.
@@ 4.1 Benchmark Experiments
In our exploration of model training dynamics, we investigate the Frequency-Principle analysis as proposed by Xu et al. [21]. This analysis method, detailed in their work, involves decomposing signals into Fourier frequencies. By employing this approach, we aim to gain deeper insights into the spectral bias phenomenon prevalent in neural network learning processes. Spectral bias refers to the tendency of neural networks to favor learning certain frequency components of input data over others, which can significantly influence model performance and generalization.
Xu et al.’s [21] Frequency-Principle analysis serves as a valuable framework for dissecting how neural networks interact with and prioritize different frequency components during training. Through this analysis, we seek to uncover patterns and tendencies in model behavior related to frequency-specific learning preferences. By understanding these dynamics, we can make informed adjustments to training strategies, network architectures, or preprocessing techniques to mitigate spectral bias effects and enhance overall model robustness and performance across various datasets and tasks.
@@ 5 Conclusion
Acknowledgment
This work was funded by NASA’s System-Wide Safety Project under NASA’s Aeronautics Mission Directorate’s Airspace Operations and Safety Program.